# Dynamic Post-Hoc Neural Ensemblers

### v1

@@ 5 Related Work
Ensembles for Tabular Data. For tabular data, ensembles are known to perform better than individual models (Sagi & Rokach, 2018; Salinas & Erickson, 2023).Therefore, ensembles are often used in real-world applications (Dong et al., 2020), to win competitions (Koren, 2009; Kaggle, 2024), and by automated machine learning (AutoML) systems as a modeling strategy (Purucker & Beel, 2023b; Purucker et al., 2023). Methods like Bagging (Breiman, 1996) or Boosting (Freund et al., 1996) are often used to boost the performance of individual models. In contrast, post-hoc ensembling (Shen et al., 2022; Purucker & Beel, 2023a) aggregates the predictions of an arbitrary set of fitted base models. Post-hoc ensembles are build by stacking (Wolpert, 1992; Van der Laan et al., 2007), ensemble selection (a.k.a. pruning) (Caruana et al., 2004; Tsoumakas et al., 2009), dynamic ensemble selection (Ko et al., 2008; Britto Jr et al., 2014), or through a systematic search for an optimal ensemble (Levesque et al., 2016; Shen et al., 2022; Poduval et al., 2024).
@@
Ensembles for Deep Learning. Ensembles of neural networks (Hansen & Salamon, 1990; Krogh & Vedelsby, 1994; Dietterich, 2000) have gained significant attention in deep learning research, both for their performance-boosting capabilities and their effectiveness in uncertainty estimation. Various strategies for building ensembles exist, with deep ensembles (Lakshminarayanan et al., 2017) being the most popular one, which involve independently training multiple initializations of the same network. Extensive empirical studies (Ovadia et al., 2019; Gustafsson et al., 2020) have shown that deep ensembles outperform other approaches for uncertainty estimation, such as Bayesian neural networks (Blundell et al., 2015; Gal & Ghahramani, 2016; Welling & Teh, 2011).
Mixture-of-Experts. Our idea of generating ensemble base model weights is closely connected to the mixture-of-experts (MoE) (Jacobs et al., 1991; Jordan & Jacobs, 1993; Shazeer et al., 2017), where one network is trained with specialized sub-modules that are activated based on the input data. Batch ensembles (Wen et al., 2020; Wenzel et al., 2020) are also closely related, as they aim to capture diverse model behaviors by simultaneously training multiple models with shared weights but different random projections. In contrast to MoE and batch ensembles, our Neural Ensemblers can ensemble any (black-box) model and are not restricted to gradient-based approaches. Ours additionally considers a regularization strategy that leads to provably more diverse ensembles.
@@
Dynamic Ensemble Selection. Our Neural Ensembler is highly related to dynamic ensemble selection. Both dynamically aggregate the predictions of base models per instance (Cavalin et al., 2013; Ko et al., 2008). Traditional dynamic ensemble selection methods aggregate the most competent base models by paring heuristics to measure competence with clustering, nearest-neighbor-based, or traditional tabular algorithms (like naive Bayes) as meta-models (Cruz et al., 2018; 2020). In contrast, we use an end-to-end trained neural network to select and weight the base models per instance.
@@ 6 Conclusions
In this work, we tackled the challenge of post-hoc ensemble selection and the associated risk of overfitting on the validation set. We introduced the Neural Ensembler, a neural network that dynamically assigns weights to base models on a per-instance basis. To reduce overfitting, we proposed a regularization technique that randomly drops base models during training, which we theoretically showed enhances ensemble diversity. Our empirical results demonstrated that Neural Ensemblers consistently form competitive ensembles across diverse data modalities, including tabular data (classification and regression), computer vision, and natural language processing. In the future, we aim to explore in-context learning (Brown et al., 2020), where a pretrained Neural Ensembler could generate base model weights at test time, using their predictions as contextual input. We discuss broader impact and limitations in Appendix B.
Reproducibility Statement
All code and datasets used in this paper are publicly available at https://github.com/machinelearningnuremberg/NeuralEnsemblers/. We provide the code for running all the experiments, including baselines on all meta-datasets. We also provide instructions for the environment setup to ensure reproducibility. We will deanonymize the link upon acceptance. We further provide all proofs in Appendix A. Moreover, we ensured reproducibility by testing our approach on many meta-datasets and across several domains, for which details can be found in Appendix D. We also compared our method to many baselines commonly used in ensembling, described in Section 4.3. Lastly, we detail additional results and extended data tables in Appendix E.
@@ Appendix A Proofs
Definition 2 (Preferred Base Model).
Consider a target variable $y\in R$ and a set of uncorrelated base models predictions $\mathcal{Z}=\{z_{m}|z_{m}\in\mathrm{R},m=1,...,M\}$. $z_{p}$ is the Preferred Base Model if it has the highest sample correlation with respect the target, i.e. $\rho_{z_{p},y}\in[0,1],\rho_{z_{p},y}>\rho_{z_{m},y},\forall z_{m}\in\mathcal{Z}/\{z_{p}\}$.
Proposition 4.
If the correlation of the preferred model $\rho_{z_{p},y}\rightarrow 1$, then $\rho_{z_{i},y}\rightarrow 0,\forall z_{i}\in\mathcal{Z}/\{z_{k}\}$, given the conditions of Definition 1.
Proof.
Let $\rho_{z_{p},z_{m}},\rho_{z_{m},y},\rho_{z_{p},y}$ be the correlations among the preferred model $z_{p}$, a non-preferred model $z_{m}\neq z_{m}$ and the target $y$, then the correlation matrix is given by,
$C=\begin{pmatrix}1&\rho_{z_{p},z_{m}}&\rho_{z_{p},y}\\ \rho_{z_{p},z_{m}}&1&\rho_{z_{m},y}\\ \rho_{z_{p},y}&\rho_{z_{m},y}&1\\ \end{pmatrix}$ (8)
The determinant of the correlation matrix is given by,
$\mathrm{Det}(C)$ $=1-\rho_{z_{p},z_{m}}^{2}-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}+2\rho_{z_{p},z_{m}}\rho_{z_{p},y}\rho_{z_{m},y}$ (9)
$\mathrm{Det}(C)$ $=1-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}$ (10)
$\mathrm{Det}(C)$ $=1-\rho_{z_{p},y}^{2}-\rho_{z_{m},y}^{2}\geq 0$ (11)
$\rho_{z_{m},y}$ $\leq\sqrt{1-\rho_{z_{p},y}^{2}},$ (12)
where we used the fact that $\rho_{z_{p},z_{m}}=0$, in Equation 10 and that the correlation matrix should be semidefinite in Equation 11, thus its determinant should hold $\mathrm{Det}(C)\geq 0$. This means that $\epsilon=\sqrt{1-\rho_{z_{p},y}^{2}}$ is an upper bound for the absolute value of the correlation $\rho_{z_{m},y}$. The limit of the upper bound $\epsilon$ as $\rho_{z_{p},y}\to 1$ is 0 and thus $\rho_{z_{m},y}=0$,
$\lim_{\rho_{z_{p},y}\to 1}$ $\epsilon=\lim_{\rho_{z_{p},y}\to 1}\sqrt{1-\rho_{z_{p},y}^{2}}=0$ (13)
$\lim_{\rho_{z_{p},y}\to 1}$ $\rho_{z_{m},y}=0.$ (14)
Since $\rho_{z_{m},y}$ is bounded by $-\epsilon\leq\rho_{z_{m},y}\leq\epsilon$.
Proposition 5 (Diversity Collapse).
If the correlation of the preferred model is $\rho_{z_{p},y}\rightarrow 1$ in an ensemble with prediction $\bar{z}=\sum_{m}\theta_{m}z_{m}$, then the ensemble diversity $\alpha\rightarrow 0$, where $\alpha:=\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$.
Proof.
Firstly, we show that under conditions of Definition 1, $\theta_{m}=\hat{\rho}_{z_{m},y}$, where $\hat{\rho}_{z_{m},y}$ is the sample correlation. Assume that we have a matrix of evaluations $X\in\mathrm{R}^{N\times M}$, with components $x_{n,m}$ indicating the prediction of the $m$-th model for the $n$-th instance. If $Y\in\mathbb{R}^{N\times 1}$ is the ground-truth, then the ensemble weights $\Theta\in\mathbb{R}^{M\times 1}$ can be computed with closed form solution of the objective $\min_{\Theta}||Y-X\Theta||^{2}$ as: $\Theta=(X^{T}X)^{-1}X^{T}Y$.
Without loss of generality, we assume that the random variables are standardized. Given that every row of $X$ have samples from uncorrelated random variables, then $X^{T}X\approx I$, denoting $I$ as the identity matrix, thus
$\Theta=X^{T}Y$ (15)
Every component of $\Theta$ can expressed as summatorias $\theta_{m}=\sum_{n}x_{n,m}y_{n}$. Note that $x_{n,m}$ is a standardized sample (i.e. mean 0, variance 1) from the random variable $z_{m}$, thus it equals the sample correlation $\theta_{m}=\sum_{n}x_{n,m}y_{n}=\hat{\rho}_{z_{m},y}$.
Now we develop the formal demonstration of Proposition 2. We note to the right side a hint to the criteria that was applied to derive every equation.
$\lim_{\rho_{z_{p},y}\to 1}\alpha=\lim_{\rho_{z_{p},y}\to 1}$ $\mathbb{E}\left[\sum_{m}\theta_{m}(z_{m}-\bar{z})^{2}\right]$ Definition of $\alpha$. (16)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\mathbb{E}\left[(z_{m}-\bar{z})^{2}\right]$ Properties of expectations. (17)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(\mathbb{E}[z_{m}-\bar{z}]^{2}+\mathbb{V}[z_{m}-\bar{z}]\right)$ Expectation of a squared variable. (18)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\mathbb{V}[z_{m}-\bar{z}]$ Means of $z_{m}$ and $\bar{z}$ are 0. (19)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(\mathbb{V}[z_{m}]-\mathbb{V}[\bar{z}]\right)$ Property of sum of variances. (20)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\mathbb{V}[\bar{z}]\right)$ Variance of $z_{m}$ is 1. (21)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\mathbb{V}\left[\sum_{m^{\prime}}\theta_{m^{\prime}}z_{m^{\prime}}\right]\right)$ Definition of $\bar{z}$. (22)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\mathbb{V}\left[z_{m^{\prime}}\right]\right)$ Properties of variance. (23)
$=\lim_{\rho_{z_{p},y}\to 1}$ $\sum_{m}\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Variance of $z_{m}$ is 1. (24)
$=\sum_{m}\lim_{\rho_{z_{p},y}\to 1}$ $\theta_{m}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Property of sum of limits. (25)
$=\sum_{m}\lim_{\rho_{z_{p},y}\to 1}$ $\theta_{m}\cdot\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\theta_{m^{\prime}}^{2}\right)$ Product-sum property of limits. (26)
$=\left[\sum_{m}\lim_{\rho_{z_{p},y}\to 1}\rho_{z_{m},y}\right]$ $\cdot\left[\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)\right]$ Given that $\theta_{m}=\hat{\rho}_{z_{m},y}\approx\rho_{z_{m},y}$. (27)
=1=1$\cdot$= 1 ⋅ $\left[\lim_{\rho_{z_{p},y}\to 1}\left(1-\sum_{m^{\prime}}\rho_{z_{m^{\prime}},y}^{2}\right)\right]$ Given Proposition 1. (28)
$=1-$ $\sum_{m^{\prime}}\lim_{\rho_{z_{p},y}\to 1}\rho_{z_{m^{\prime}},y}^{2}$ Properties of limits. (29)
$\lim_{\rho_{z_{p},y}\to 1}$ $\alpha=0$ Given Proposition 1. (30)
Lemma 1.
When using dropout, it still holds that $\theta_{m}=\hat{\rho}_{z_{m},y}.$
Proof.
$\min_{\Theta}\mathbb{E}_{R\sim\mathrm{Ber(\gamma)}}||Y-(R*X)\Theta||^{2}$ (31)
$=\min_{\Theta}||Y-\gamma X\Theta||^{2}+\gamma(1-\gamma)||\Gamma\Theta||^{2},$ (32)
where $\Gamma=(\mathrm{diag(X^{T}X)})^{1/2}$. If we use the uncorrelation of the base models, we obtain:
$=\min_{\Theta}||Y-\gamma X\Theta||^{2}+\gamma(1-\gamma)||\Theta||^{2}$ (33)
$=\min_{\Theta}||Y-\tilde{X}\Theta||^{2}+\gamma(1-\gamma)||\Theta||^{2},$ (34)
which is equivalent to ridge regression, with closed form solution:
$\Theta=$ $\left({\tilde{X}}^{T}{\tilde{X}}+\gamma(1-\gamma)I\right)^{-1}{\tilde{X}}^{T}Y$ (35)
$=$ $\left([\gamma^{2}+\gamma(1-\gamma)]I\right)^{-1}{\tilde{X}}^{T}Y$ (36)
$=$ $\left(\gamma I\right)^{-1}{\tilde{X}}^{T}Y$ (37)
$=$ $\frac{1}{\gamma}{\tilde{X}}^{T}Y$ (38)
$=$ $X^{T}Y.$ (39)
Proposition 6 (Avoiding Diversity Collapse).
@@
We follow a similar procedure as for Proposition 2, by considering $\bar{z}=\sum_{m}r_{m}\ \theta_{m}z_{m}$. We demonstrate that $\mathbb{V}(r\cdot z_{m})=\gamma$, given that $r\sim\mathrm{Bernoulli}(\gamma)$.
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\mathbb{V}(r_{m})\cdot\mathbb{V}(z_{m})+\mathbb{V}(z_{m})\cdot\mathbb{E}(r_{m})^{2}+\mathbb{V}(r_{m})\cdot\mathbb{E}(z_{m})^{2}$ (40)
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\mathbb{V}(r_{m})+\mathbb{E}(r_{m})^{2}$ (41)
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\gamma(1-\gamma)+\gamma^{2}$ (42)
$\mathbb{V}(r_{m}\cdot z_{m})$ $=\gamma.$ (43)
Then, we evaluate the variance of the ensemble using DropOut $\mathbb{V}(\bar{z})$:
$\mathbb{V}(\bar{z})=\mathbb{V}{\left(\sum_{m}r_{m}\cdot\theta_{m}\cdot z_{m}\right)}$ (44)
$\mathbb{V}(\bar{z})=\sum_{m}\mathbb{V}{\left(r_{m}\cdot\theta_{m}\cdot z_{m}\right)}$ (45)
$\mathbb{V}(\bar{z})=\sum_{m}\theta^{2}_{m}\mathbb{V}{\left(r_{m}\cdot z_{m}\right)}$ (46)
$\mathbb{V}(\bar{z})=\gamma\sum_{m}\theta^{2}_{m}.$ (47)
Applying Equation 47 into Equation 22, we obtain:
$\lim_{\rho_{z_{p},y}\to 1}\alpha$ $=\lim_{\rho_{z_{p},y}\to 1}\mathbb{E}\left[\sum_{m}\gamma_{m}\cdot\theta_{m}(z_{m}-\bar{z})^{2}\right]$ (48)
@@
Appendix B Limitations and Broader Impact.
@@
While our proposed method offers several advantages for post-hoc ensemble selection, it is important to recognize its limitations. Unlike simpler ensembling heuristics, our approach requires tuning multiple training and architectural hyperparameters. Although we employed a fixed set of hyperparameters across all modalities and tasks in our experiments, this robustness may not generalize to all new tasks. In such cases, hyperparameter optimization may be necessary to achieve optimal performance. However, this could also enhance the results presented in this paper. Our approach is highly versatile and can be seamlessly integrated into a wide variety of ensemble-based learning systems, significantly enhancing their predictive capabilities. Because our method is agnostic to both the task and modality, we do not expect any inherent negative societal impacts. Instead, its effects will largely depend on how it is applied within different contexts and domains, making its societal implications contingent on the specific use case.
Appendix C Related Work Addendum

### v2

@@ 5  Related Work
Ensembles for Tabular Data. For tabular data, ensembles are known to perform better than individual models (Sagi and Rokach,, 2018; Salinas and Erickson,, 2023).Therefore, ensembles are often used in real-world applications (Dong et al.,, 2020), to win competitions (Koren,, 2009; Kaggle,, 2024), and by automated machine learning (AutoML) systems as a modeling strategy (Purucker and Beel, 2023b, ; Purucker et al.,, 2023). Methods like Bagging (Breiman,, 1996) or Boosting (Freund et al.,, 1996) are often used to boost the performance of individual models. In contrast, post-hoc ensembling (Shen et al.,, 2022; Purucker and Beel, 2023a, ) aggregates the predictions of an arbitrary set of fitted base models. Post-hoc ensembles can be built by stacking (Wolpert,, 1992) ensemble selection (a.k.a. pruning) (Caruana et al.,, 2004) , or through a systematic search for an optimal ensemble (Levesque et al.,, 2016).
@@
Ensembles for Deep Learning. Ensembles of neural networks (Hansen and Salamon,, 1990; Krogh and Vedelsby,, 1994; Dietterich,, 2000) have gained significant attention in deep learning research, both for their performance-boosting capabilities and their effectiveness in uncertainty estimation. Various strategies for building ensembles exist, with deep ensembles (Lakshminarayanan et al.,, 2017) being the most popular one, which involves independently training multiple initializations of the same network. Extensive empirical studies (Ovadia et al.,, 2019; Gustafsson et al.,, 2020) have shown that deep ensembles outperform other approaches for uncertainty estimation, such as Bayesian neural networks (Blundell et al.,, 2015; Gal and Ghahramani,, 2016; Welling and Teh,, 2011).
@@
Dynamic Ensemble Selection. Our Neural Ensembler is highly related to dynamic ensemble selection. Both dynamically aggregate the predictions of base models per instance (Cavalin et al.,, 2013; Ko et al.,, 2008). Traditional dynamic ensemble selection methods aggregate the most competent base models by paring heuristics to measure competence with clustering, nearest-neighbor-based, or traditional tabular algorithms (like naive Bayes) as meta-models (Cruz et al.,, 2018, 2020). In contrast, we use an end-to-end trained neural network to select and weight the base models per instance.
Mixture-of-Experts. Our idea of generating ensemble base model weights is closely connected to the mixture-of-experts (MoE) (Jacobs et al.,, 1991; Jordan and Jacobs,, 1993; Shazeer et al.,, 2017), where one network is trained with specialized sub-modules that are activated based on the input data. Alternatively, we could include a layer, aggregating predictions by encouraging diversity (Zhang et al.,, 2020). In contrast to these approaches, our Neural Ensemblers can ensemble any (black-box) model and are not restricted to gradient-based approaches.
@@ 6  Conclusions
In this work, we tackled the challenge of post-hoc ensemble selection and the associated risk of overfitting on the validation set. We introduced the Neural Ensembler, a neural network that dynamically assigns weights to base models on a per-instance basis. To reduce overfitting, we proposed a regularization technique that randomly drops base models during training, which we theoretically showed enhances ensemble diversity. Our empirical results demonstrated that Neural Ensemblers consistently form competitive ensembles across diverse data modalities, including tabular data (classification and regression), computer vision, and natural language processing.
@@ Acknowledgments
We want to summarize here all the appendix sections:
  • Section A presents the proofs of propositions in the main paper.
  • Section B discusses the limitations of our proposed method and its broader impacts.
  • Section D further details research similar to our work.
  • Section O explains the process of gathering and preparing the FTC and Scikit-learn Pipelines metadatasets used in our experiments.
  • Section F includes supplementary tables and figures, such as average ranks and detailed performance metrics, that support and expand upon the main experimental results reported in the paper.
  • Section H includes he computational cost associated with our method compared to baseline approaches, including runtime evaluations and discussions on efficiency.
  • Section I includes the results of a proof-of-concept experiment using overparameterized base models (e.g., 10th-degree polynomials), demonstrating the effectiveness of our Neural Ensembler even when base models have high capacity.
  • Section G includes the Critical Difference diagrams corresponding to our main results, illustrating the statistical significance of performance differences among methods and how to interpret these diagrams.
  • Section J explores the impact of using different amounts of validation data to train the Neural Ensembler, assessing its sample efficiency and how performance scales with varying data sizes.
  • Section K explores the effect of merging the training and validation datasets on the performance of both base models and ensemblers.
  • Section L explores an alternative formulation of the Neural Ensembler that operates on the original input space rather than on the predictions of the base model, including experimental results and discussions of its effectiveness.
  • In section M, we study whether the performance of the models changes when using base models found by Bayesian Optimization or randomly.
@@ Appendix A Proofs
Proposition A.1 (Diversity Lower Bound).
@@
[…]
Applying Equation 15 into Equation LABEL:proof2:l7, we obtain:
[…]
@@
Appendix B Limitations, Broader Impact and Future Work
@@
While our proposed method offers several advantages for post-hoc ensemble selection, it is important to recognize its limitations. Unlike simpler ensembling heuristics, our approach requires tuning multiple training and architectural hyperparameters. Although we employed a fixed set of hyperparameters across all modalities and tasks in our experiments, this robustness may not generalize to all new tasks. In such cases, hyperparameter optimization may be necessary to achieve optimal performance. However, this could also enhance the results presented in this paper. Additionally, some bayesian approaches (McEwen et al.,, 2021) could further increase the robustness to the size of the validation data set.
Our approach is highly versatile and can be seamlessly integrated into a wide variety of ensemble-based learning systems, significantly enhancing their predictive capabilities. Because our method is agnostic to both the task and modality, we do not expect any inherent negative societal impacts. Instead, its effects will largely depend on how it is applied within different contexts and domains, making its societal implications contingent on the specific use case. In the future, we aim to explore in-context learning (Brown et al.,, 2020), where a pretrained Neural Ensembler could generate base model weights at test time, using their predictions as contextual input. We discuss broader impact and limitations in Appendix B.
Appendix C Space complexity of Neural Ensemblers
[…]
Appendix D Related Work Addendum